# New Online Communities: Graph Deep Learning on Anonymous Voting Networks to Identify Sybils in Polycentric Governance

### v6

@@ 5.3.2 Graph Time Signal Analysis
In the pursuit of elucidating the complex temporal dynamics of DAOs, we engineer and embed spectral time series data, use wavelet transforms with unsupervised clustering techniques to reduce the high dimensional signal and visualize with UMAP. This multi-resolution analysis is particularly useful for highlighting non-linear patterns and dynamics obscured in raw time series data, making wavelets a suitable preprocessing step for uncovering spectral time embeddings.
DRAFT: INSERT DIAGRAMS
Following the wavelet transformation, clustering algorithms are used to group vectors based on similarity of their spectral and temporal features, revealing distinct patterns or behaviors within the dataset. Clustering is important for understanding the inherent structure and relationships within the data, providing a basis for further analysis and visualization. Finally, UMAP is used to project the high-dimensional clusters into a lower-dimensional space, preserving the essential topological structure of the data. This dimensionality reduction enables a visual exploration of the clustered spectral time embeddings and underscores the relationships and distinctions between identified patterns. The amalgamation of wavelet-based feature extraction, clustering, and UMAP visualization presents an experimental framework for analyzing time series data, helping unearth and visualize complex spectral time dynamics that would otherwise remain concealed.
DRAFT: INSERT DIAGRAMS
@@ 5.3.3 Wallet Fingerprint Signal Analysis
While we cannot directly validate Sybils by tracing origin chain transactions, we can use historical wallet balances to “fingerprint” voters. Using the snapshot.org API, we constructed a graph of voters for ten DAOs of varying sizes and augmented each voter with unique wallet token balances at the time of the vote. We used the Covalent Unified API to collect token tickers and their associated historical balances and attached these ‘coins’ to each voter. From Figures 14 and 14, we see many unique tokens in voter wallets and observe that voters share wallet token sets, with a long tail of shared, distinct voter sets. If we then dot plot the token sets and colormap their normalized values, we see clear evidence of extractable information for each wallet defined by the uniqueness of wallet token sets (Figure 15).
@@
The resulting clusters can be evaluated against the true labels with a confusion matrix (Figure 22). Additionally, we computed cluster evaluation metrics (see 1).
@@ 5.3.4 Graph Convolutional Neural Network (GCNN)
For our method, the first few steps are relatively straightforward. First, construct a voting graph where known (labelled) voters and re-used wallets are combined into single nodes with multiple edges to proposal nodes (we use a multi- or hyper-graph to accommodate multiple edges from each node). Second, we engineer features to learn higher dimensional vector embeddings on all nodes (this is the encoder). We use a simple reconstruction loss (Mean Squared Error or MSE) to capture node features for training the GCNN (we do not use a more conventional classification training loss like cross-entropy because we do not know the Sybil classes a priori). Third, we use the node embeddings to conduct a k-means vector clustering on unknown nodes (this is the decoder). We then propagate the most common label associated with each cluster to relabel the actual, lower dimensional nodes to form the Sybil cluster (this is also a type of MaxPooling layer for dimensionality reduction). Finally, we take the relabelled similarity graph and combine nodes with their labels in each cluster to form a higher dimensional, but reduced-size graph (upscaling). The resulting graph is an inductive prediction of Sybil nodes.
@@
The deep learning method was developed with the Python libraries NetworkX and RAPIDS AI to construct the graphs, the PyTorch Geometric library for the GCNN, and the Facebook AI Similarity Search (FAISS) for the fast embeddings clustering. Whenever possible, GPU accelerated libraries were used, although speed and memory tradeoffs quickly become a practical data analysis challenge when dealing with large graphs requiring several hundred gigabytes of memory for large matrix multiplications during learning. This is known as the graph expansion problem and can be controlled by neighborhood sampling or graph partitioning (batch normalization). As an inductive method, decisions to use a particular library, feature, or machine learning algorithm were made pragmatically and tested against the dataset, but as discussed below, the GCNN encoder-decoder approach offers some distinct conceptual advantages over other approaches.
@@
We develop features and construct a typical train/test/validation split for the forward pass of our GCNN. The first layer of the forward pass is fully connected, capturing node features and incorporating voting edges with their ‘power’ attributes. The second layer integrates edge and node features with a sequential multi-layer perceptron (MLP) using ReLu activation, which is then fed into a convolutional layer with mean feature aggregation. The third layer processes temporal sequences engineered for a Long Short-Term Memory (LSTM) layer. Finally, the fourth layer uses a multi-headed Graph Attention Network (GAT) to better learn and search edge connectivity. Due to the final attention layer, this encoder can be considered a transformer with residual connections. (Due to dataset availability limitations, the wallet fingerprint signal was not used.)
@@ 5.3.5 Fast Vector Embeddings Search using FAISS
Once we have learned higher dimensional voter and voting representations for all (known and unknown) nodes, we can cluster similar vector embeddings (an encoding task), which we initially learned using the GCNN. In deciding on our method, we also considered other similarity metrics like Euclidian distance, cosine similarity, and Jaccard similarity, but computing many-to-many similarity searches on large matrixes is very slow $O(n^{2}\cdot d)$. Thankfully, the Facebook AI Similarity Search (FAISS) Python library provides extremely fast and easily scaled search and clustering options (including GPU processing) for vector embeddings. FAISS offers several search index options for improving efficiency, including approximate search methods that can scale to billions of elements. For our purposes, the L2 index was quick enough and provides exact results (for much larger datasets, FAISS provides approximate methods that are substantially quicker). We use the included k-means clustering method to find Sybil clusters, which uses centroids in the high dimensional embeddings to cluster and return neighbours. We normalize the clusters by dropping any remaining singletons and large clusters (mean + 1 standard deviation).
@@
We then inspect each cluster and propagate the most common label across the cluster to construct a “similarity” graph (downsampling). Finally, we construct a reduced “clustered” graph by combing clustered nodes (and reconstructing edges) (upsampling).
@@
In formal notation, the complete method:
@@ 6   Results
We have conducted numerous graph analysis experiments on DAO governance voting with a range of graph sizes, learnable features, specialized convolutional layers, and training and optimization algorithms. We also tested different clustering approaches including DBMSCAN and other unsupervised methods; we deployed supervised classification methods with the labelled subgraph of true (Ethereum Name Service) identities; and we considered the problem from different graph analysis viewpoints (such as considering Sybil identification as a problem of link prediction or as a community detection problem). While many of these methods initially looked promising, many failed for a variety of reasons; most commonly, approaches cannot rely on (k$>$1) edge connectivity, which is precisely what provides anonymity.
@@
Graph convolutional neural networks offer several advantages for identifying Sybils. First, network topology is informative but difficult to learn using traditional methods. This is especially true for very large graphs. However, since convolutional networks aggregate information from nearby nodes, and therefore inductively bias these relationships, learning is simplified (the graph signal is downsampled), and since parameters are shared across convolutional layers (lowering memory requirements), large graphs are less difficult to learn. Moreover, our encoder-decoder approach takes cues from Natural Language Processing (NLP), which must accommodate a large number of token inputs (precluding deep, fully connected networks). Our use of a multi-head graph attention layer produces node embeddings with n-deep neighborhood information (this is also known as a transformer), thereby increasing network depth but with tractable sparsity. Similarly, the Long-Short Term Memory (LSTM) algorithm used to learn time series data can be considered an earlier version of the transformer model. More broadly, transformer models, although still rarely applied to graph problems, offer significant untapped potential for network deanonymization and have already proven useful for image processing and other discrete structural problems.
@@
Finally, we used the trained embeddings to generatively produce voting samples, which we stimulated with varying Sybil signals. This voting simulation accurately reflects the temporal dynamics and evolution of online communities by generating samples from high dimensional embeddings representing specific voting behaviours. The introduction of Sybil signals reveals potential sources of resource congestion and over extraction.
DRAFT: INSERT DIAGRAMS
@@ 7   Future Research
To improve and further validate our inductive method of Sybil identification we need benchmarks from other online communities. Moreover, more sophisticated and further development of machine learning features would enhance the method. Finally, despite the challenges required to ‘follow the money,’ using this direct method to identify Sybil chains is the most reliable signal we have, and if conducted at scale the method would provide an invaluable graph signal for supervised learning.
In addition to refining current methods, it is worth exploring other deep learning techniques. To adapt the method for detecting voting spam, we could instead detect anomalies within the node embeddings. Instead of identifying similar nodes, searching for outliers might indicate spam activity (perhaps this might be exploited using transfer learning). Or, since the GCNN trains an embeddings model, generative tasks like link prediction or economic simulation might be possible.
Finally, deep learning methods can be exploited by malicious users. Making deep learning methods resilient to attack is an active area of research and a practical requirement for DAOs, which must protect their digital commons from theft and inappropriate extraction. Examples of attacks against ML models include adversarial attacks that manipulate input data to deceive models, model inversion attacks that reveal sensitive information, and poisoning attacks that corrupt training data. Developing robust models that can detect and resist these attacks is crucial and should be part of any production environment. Additionally, understanding and mitigating the ethical implications of AI and ML in DAO governance, such as bias in decision-making processes, must be a priority. These challenges open up diverse research avenues, ranging from developing new algorithms for anomaly detection to ethical frameworks for AI deployment in decentralized governance structures.
@@ 8   Conclusion
The Internet’s transformative impact on communication costs has laid the groundwork for a similar paradigm shift in governance technologies. By drastically reducing the costs associated with dispute resolution and collaboration, these technologies have the potential to alter the landscape of governance. This shift is frustrated by anonymous accounts called Sybils, but by deploying a deep learning method we augmented network transparency to minimize the costs of monitoring and enforcement within governance structures. Without impacting personal privacy and without relying on unproven identity systems, our approach modulates governance behaviours at scale to prevent Sybil attacks, potentially heralding a new era of efficiency and effectiveness in governance systems.
The concept of Digital Common Pool Resources (DCPRs) and polycentric governance emerges as a promising avenue for future research in this context. This governance model is particularly relevant in addressing the perennial challenges associated with managing digital assets in a commons, such as congestion, overuse, pollution, and destruction. Traditional approaches often struggle with enforcing limits on resource utilization, but polycentric governance offers a novel framework that could be more effective in managing these resources. With polycentric governance, participants in the digital commons develop context-specific rules to prevent infringement and establish consequences for those who exploit the commons. By leveraging decentralized decision-making and consensus mechanisms, these regimes hold the potential for more sustainable and equitable resource management.

### v7

@@ 5.3.2 Graph Time Signal Analysis
In the pursuit of elucidating the complex temporal dynamics of DAOs, we engineer and embed spectral time series data, use wavelet transforms with unsupervised clustering techniques to reduce the high dimensional signal. This multi-resolution analysis is particularly useful for highlighting non-linear patterns and dynamics obscured in raw time series data, making wavelets a suitable preprocessing step for uncovering spectral time embeddings.
Following the wavelet transformation, clustering algorithms are used to group vectors based on similarity of their spectral and temporal features, revealing distinct patterns or behaviors within the dataset. Clustering is important for understanding the inherent structure and relationships within the data, providing a basis for further analysis and visualization. Finally, reducing the high-dimensional clusters into a lower-dimensional space preserves the essential topological structure of the data. This dimensionality reduction enables a visual exploration of the clustered spectral time embeddings and underscores the relationships and distinctions between identified patterns. The amalgamation of wavelet-based feature extraction, clustering, and visualization presents an experimental framework for analyzing time series data, helping unearth and visualize complex spectral time dynamics that would otherwise remain concealed.
@@ 5.3.3 Wallet Fingerprint Signal Analysis
While we cannot directly validate Sybils by tracing origin chain transactions, we can use historical wallet balances to “fingerprint” voters. Using the snapshot.org API, we constructed a graph of voters for seven DAOs of varying sizes and augmented each voter with unique wallet token balances at the time of the vote. We used the Covalent Unified API to collect token tickers and their associated historical balances and attached these ‘coins’ to each voter. From Figures 14 and 14, we see many unique tokens in voter wallets and observe that voters share wallet token sets, with a long tail of shared, distinct voter sets. If we then dot plot the token sets and colormap their normalized values, we see clear evidence of extractable information for each wallet defined by the uniqueness of wallet token sets (Figure 15).
Figure 13: Histogram of Unique Wallet Tokens.
@@
Figure 17: UMAP Projection of Wallet Token Fingerprints.
@@
The resulting clusters can be evaluated against the true labels with a confusion matrix (Figure 22). Additionally, we computed cluster evaluation metrics (see Table 1).
@@ 5.3.4 Graph Convolutional Neural Network (GCNN)
For our method, the first few steps are relatively straightforward. First, construct a voting graph where known (labelled) and unknown voters are combined into single nodes with multiple edges to proposal nodes (we use a multi- or hyper-graph to accommodate multiple edges from each node). Second, we engineer features to learn higher dimensional vector embeddings on all nodes (this is the encoder). We use a simple reconstruction loss (Mean Squared Error or MSE) to capture node features for training the GCNN (we do not use a more conventional classification training loss like cross-entropy because we do not know the Sybil classes a priori). Third, we use the node embeddings to conduct a k-means vector clustering on unknown nodes (this is the decoder). We then propagate the most common label associated with each cluster to relabel the actual, lower dimensional nodes to form the Sybil cluster (this is also a type of MaxPooling layer for dimensionality reduction). Finally, we take the relabelled similarity graph and combine nodes with their labels in each cluster to form a higher dimensional, but reduced-size graph (upscaling). The resulting graph is an inductive prediction of Sybil nodes.
@@
The deep learning method was developed with the Python libraries NetworkX and RAPIDS AI to construct the graphs, the PyTorch Geometric library for the GCNN, and the Facebook AI Similarity Search (FAISS) for the fast embeddings clustering. Whenever possible, GPU accelerated libraries were used, although speed and memory tradeoffs quickly become a practical data analysis challenge when dealing with large graphs requiring several hundred gigabytes of memory for large matrix multiplications during learning. This is known as the graph expansion problem and can be controlled by neighborhood sampling or graph partitioning (in our case, batch normalization). As an inductive method, decisions to use a particular library, feature, or machine learning algorithm were made pragmatically and tested against the dataset, but as discussed below, the GCNN autoencoder approach offers some distinct conceptual advantages over other approaches.
@@
We develop features and construct a typical train/test/validation split for the forward pass of our GCNN. The first layer of the forward pass is fully connected, capturing node features and incorporating voting edges with their ‘power’ attributes. The second layer integrates edge and node features with a sequential multi-layer perceptron (MLP) using ReLu activation, which is then fed into a convolutional layer with mean feature aggregation. The third layer processes temporal sequences engineered for a Long Short-Term Memory (LSTM) layer. Finally, the fourth layer uses a multi-headed Graph Attention Network (GAT) to better learn and search edge connectivity. Due to the final attention layer, this encoder can be considered a transformer with residual connections. (Note that due to dataset availability limitations, the wallet fingerprint signal was not used in this final algorithm.)
@@ 5.3.5 Fast Vector Embeddings Search using FAISS
Once we have learned higher dimensional voter and voting representations for all (known and unknown) nodes, we can cluster similar vector embeddings (an encoding task), which we initially learned using the GCNN. In deciding on our method, we also considered other similarity metrics like Euclidian distance, cosine similarity, and Jaccard similarity, but computing many-to-many similarity searches on large matrixes is very slow; $O(n^{2}\cdot d)$. Thankfully, the Facebook AI Similarity Search (FAISS) Python library provides extremely fast and easily scaled search and clustering options (including GPU processing) for vector embeddings. FAISS offers several search index options for improving efficiency, including approximate search methods that can scale to billions of elements. For our purposes, the L2 index was quick enough and provides exact results (for much larger datasets, FAISS provides approximate methods that are substantially quicker). We use the included k-means clustering method to find Sybil clusters, which uses centroids in the high dimensional embeddings to cluster and return neighbours. We normalize the clusters by dropping any remaining singletons and large clusters (mean + 1 standard deviation).
@@
We then inspect each cluster and propagate the most common label across the cluster to construct a “similarity” graph (downsampling). Finally, we construct a reduced “clustered” graph by combing clustered nodes and reconstructing edges (upsampling).
@@
In formal notation, the complete experimental Sybil identification method:
@@ 6   Results
We have conducted numerous graph analysis experiments on DAO governance voting with a range of graph sizes, learnable features, specialized convolutional layers, and training and optimization algorithms. We also tested different clustering approaches including DBSCAN and other unsupervised methods; we deployed supervised classification methods with the labelled subgraph of true (Ethereum Name Service) identities; and we considered the problem from different graph analysis viewpoints (such as considering Sybil identification as a problem of link prediction or as a community detection problem). While many of these methods initially looked promising, most failed for a variety of reasons; most commonly, approaches cannot rely on (k$>$1) edge connectivity, which is precisely what provides voter anonymity.
@@
Graph convolutional neural networks offer several advantages for identifying Sybils. First, network topology is informative but difficult to learn using traditional machine learning methods. This is especially true for very large graphs. However, since convolutional networks aggregate information from nearby nodes, and therefore inductively bias these relationships, learning is simplified (the graph signal is downsampled), and since parameters are shared across convolutional layers (lowering memory requirements), large graphs are less difficult to learn. Moreover, our encoder-decoder approach takes cues from Natural Language Processing (NLP), which must accommodate a large number of token inputs (precluding deep, fully connected networks). Our use of a multi-head graph attention layer produces node embeddings with n-deep neighborhood information (this is also known as a transformer), thereby increasing network depth but with tractable sparsity. Similarly, the Long-Short Term Memory (LSTM) algorithm used to learn time series data can be considered an earlier version of the transformer model. More broadly, transformer models, although still rarely applied to graph problems, offer significant untapped potential for network deanonymization and have already proven useful for image processing and other discrete structural problems.
@@ 7   Future Research
To improve and further validate our inductive method of Sybil identification we need benchmarks from other online communities. Moreover, more sophisticated and further development of machine learning features would enhance the method. Finally, despite the challenges required to ‘follow the money,’ using this direct method to identify Sybil chains is the most reliable signal we have, and if conducted at scale the method would provide an invaluable graph signal for additional supervised learning.
We could also use the trained embeddings to generatively produce voting samples, which can be stimulated with varying Sybil signals. By generating a sampled voting signal, we can create an accurate simulation to reflect the temporal dynamics and evolution of online communities. The insight for simulation is that by generating samples from high dimensional embeddings we can represent specific voting behaviours. Further, the introduction of Sybil signals reveals potential sources of resource congestion and over-extraction.
In addition to refining current methods, it is worth exploring other deep learning techniques. To adapt the method for detecting voting spam, we could instead detect anomalies within the embeddings. Instead of identifying similar nodes, searching for outliers might indicate spam activity (perhaps this might be exploited using transfer learning). Or, since the GCNN trains an embeddings model, generative tasks like link prediction or more complex temporal-economic simulation might be possible.
Finally, machine learning methods can be exploited by malicious users. Making machine learning methods resilient to attack is an active area of research and a practical requirement for DAOs, which must protect their digital commons from theft and inappropriate extraction. Examples of attacks against ML models include adversarial attacks that manipulate input data to deceive models, model inversion attacks that reveal sensitive information, and poisoning attacks that corrupt training data. Developing robust models that can detect and resist these attacks is crucial and should be part of any production environment (see [10]). Additionally, understanding and mitigating the ethical implications of AI and ML in DAO governance, such as bias in decision-making processes, must be a priority. These challenges open up diverse research avenues, ranging from developing new algorithms for anomaly detection to ethical frameworks for AI deployment in decentralized governance structures.
@@ 8   Conclusion
The Internet’s transformative impact on communication costs has laid the groundwork for a similar paradigm shift in governance technologies. By drastically reducing the costs associated with dispute resolution and collaboration, these technologies have the potential to alter the landscape of online life. This shift is frustrated by anonymous accounts called Sybils, but by deploying a deep learning method we augmented network transparency to minimize the costs of monitoring and enforcement within governance structures. Without impacting personal privacy and without relying on unproven identity systems, our approach modulates governance behaviours at scale to prevent Sybil attacks, potentially heralding a new era of efficiency and effectiveness in governance systems.
The concept of Digital Common Pool Resources (DCPRs) and polycentric governance emerges as a promising avenue for future research in this context. This governance model is particularly relevant in addressing the perennial challenges associated with managing digital assets in a commons, such as congestion, overuse, pollution, and theft. Traditional approaches often struggle with enforcing limits on resource utilization, but polycentric governance offers a novel framework that could be more effective in managing these resources. With polycentric governance, participants in the digital commons develop context-specific rules to prevent infringement and establish consequences for those who exploit the commons. By leveraging decentralized decision-making and consensus mechanisms, these regimes hold the potential for more sustainable and equitable digital resource management.